OpenStudy (anonymous):

If a sub n = 6 + (n - 1)5, then a sub 7 =

OpenStudy (anonymous):

a(7) = 6+(7-1)*5 a(7) = 6+(6*5), 6*5 = 30 a(7) = 6+30 = 36